ah42632:

What is the value of x? A. 5 B. 2.5 C. 7.5 D. 10

ah42632:

here is the picture of the triangle

ah42632:

darkknight:

similar triangles we can say that \[x/(x+5)=(x-2)/(x+1)\] because if we look at the 2 similar triangles, there is a common ratio between them a side of the long triangle divided by a corresponding side from the short = another side from the long triangle divided by the corresponding side of that other side

darkknight:

the next step here is to isolate x from the denominator, \[x/(x+5)=(x-2)/(x+1)\] \[x(x+1)=(x+5)(x-2)\]

darkknight:

with me so far?

ah42632:

yep

darkknight:

\[x(x+1)=(x+5)(x-2)\] \[x^2+x = x^2+5x-2x-10\] \[x^2+x = x^2+3x-10\] subtract both sides by x^2 think you can go from here?

ah42632:

yeah. thx so much

darkknight:

np